FEL Upgrade Project Weekly Brief
October 23-27, 2000


We survived an intensive review by the ONR/DOE semi-annual project review committee that grilled us from  08:00 to 18:00 on Monday, Oct. 23.  According to our ONR contract monitor, John Albertine: “the bottomline-it looks like the project is going well”.  Details will follow when the committee report is filed.  Thanks to all the project team and the external reviewers for their effort in preparing for and participating in the review.


The project monthly report for September was completed in preparation for this week’s semi-annual review.  It will be distributed to ONR and DOE today.  We have one more near term review.  Next week, FEL project
management will report to an external review team that reviews Jefferson Lab’s management on a biannual basis as a part of SURA’s performance metrics for the DOE contract.

WBS 3 (Beam Physics)

Following the review, activities centered on an examination on performance limits in energy recovering linacs, with particular attention on linac dynamic range inasmuch as this parameter is 5 times larger in the Upgrade than in the demo (10 MeV:200 MeV vs. 10 MeV:40 MeV).

It was determined that a very large injected:final:recovered energy ratio could be supported through the use appropriate external focussing in the linac.

WBS 4 (Injector)

Installed the 2nd N2 implanted electrode into the FET system and it is baking.We are looking for repeatability in the performance of the implanted electrode which held 30MV/m.

Fuses were replaced for the variac in the HVPS for cavity 4 of the Injector.  The system loop gain was also lowered to stop oscillations.

The requisition for the Gun's HVPS is still in Purchasing.  The Purchase Order has not yet been placed.


Received weld coupons from PHPK for qualification of helium vessel heads and a sample vessel head. Plan to visit PHPK next week to inspect 1st article helium vessel prior to shipment to Jefferson Lab.

WBS 6 (RF)

Debug of the interlocks for RF system for zone 4 continued this week.

WBS 8  (I & C)

Four candidates were interviewed for the FEL I&C coordinator position, each had varied strengths; discussions will continue next week. Commissioning on the picomotor multiplexer continued this week during "down" opportunities.  Sub-systems needed for the optical BPMs are being completed and installed.  The detector preamps are progressing.

WBS 9 (Transport)

 We continued generating the GW 3D layout making considerable progress in integrating leads, bus, cover, hoses and bolting. The magnetic model of the GX ran but yielded one third the expected field. We are debugging
the imput file.
 Next Priorities:
(1) Determine effective length and K1 from existing baseline-GX model
(2)Build brand new 3D model of modified-GX

QX (3.125" Quad) Magnets
 Prototype Fabrication: The vendor identified a potential source of the offset and will take another cut on the same part.  This will be CMM'd once again to check for conformity.  The modified part is expected early next week.
Budgetary Quote: We will send the coil and quadrant packages out separately to additional vendors to get a better assessment of the industrial price.
Measurement Probe: Probe material was narrowed down to two choices:
Pyrex® glass (as previous probes) or a reinforced plastic compound (Ultem PEI 2300.)  We will pursue getting samples of both materials to test machinability and bonding.  The Jlab Shop offered to help test samples but their workload will not permit fabricating the final probe.

 The distribution of the drafting /design load to be done using outside Engineering Services was firmed up to all remaining dipoles and their chambers.

WBS 10 (Wiggler)

An engineer was assigned to monitor the fabrication of the dispersion section.  The steel for the dispersion section was cut by the machine shop and sent to PECO.  The decision was made to measure the optical klystron in house using a Hall probe.  We now need to decide which Hall probe to use and how to move the probe through the optical klystron.  Two options exist for each question.  We are researching the two issues and should be able to come up with a plan next week.
Other items:
Assembly: Awaiting modified clamp hardware being fabricated.
Drawings:  details of the manifold ends and modifications of the straight parts were signed off.
Manifold brackets and buss bar mounting plates are next.

WBS 11 (Optics)

This week, our upgrade efforts concentrated on the following areas:

The other O-BPM's mechanical assembly was installed in the Optics Control Room (OCR).  We are adding the exterior optical components and hope to begin testing them soon.  We have run more tests of the beam
position monitoring functionality provided by the PyroCAM IR camera in the OCR and the ModeMaster, located at the end-of-line (EOL) diagnostic hutch.  These have proven very useful while running in cw mode.

The mirror test stand vacuum enclosure prints are done, but we need to confirm several of the internal component placements before sending the package out for bid.

As for operations efforts, we've learned that the all-reflective condensing telescope in the OCR efficiently transports harmonics to the autocorrelator, and prevented us from obtaining data.  Filters will be procured to remedy this.  The 50:50 beamsplitter, which worked well at 5 microns, is more like 25:75 at 6 microns.  It will be replaced as well.


This week the FEL was operated for the following users: Mike Kelley for polymer ablation studies; B. Holloway et al. for carbon nanotube studies; and (3) the Duke Medical School group for corneal ablation studies.

Privacy and Security Notice